OpenStudy (anonymous):

I need help with 1+42^13=

OpenStudy (anonymous):

Is this estimating?

OpenStudy (anonymous):

1265437718438866624513?

OpenStudy (anonymous):

well here is the original problem

OpenStudy (anonymous):

Let g(x)= -x^(-2) + 6x^14 + pi^7. Find g'(0.5).

OpenStudy (anonymous):

and I got to..

OpenStudy (anonymous):

Yay you go me to level 6 woot!

OpenStudy (amistre64):

clean up the replies please ....

OpenStudy (anonymous):

g(x)= 2(0.5)^-3+84(0.5)^13

OpenStudy (anonymous):

Who was that redirrected towards ?

OpenStudy (amistre64):

looks like something for a calculator :)

OpenStudy (amistre64):

yes, it was.

OpenStudy (anonymous):

So you get a derived function: 2x + 84x^13 + 7pi And now: 2(0.5) + 42 * 0.5^13 + 7pi = 22.99628... Ce'nest pa?

OpenStudy (anonymous):

Yes it sure does Amistre64 my caculator messed up a bit :P i thought i was messed up

OpenStudy (anonymous):

1+42^13

OpenStudy (amistre64):

:) i tend to use the wolf for those things .... makes life simpler on the cuticles

OpenStudy (anonymous):

what is the owlf?

OpenStudy (anonymous):

wolf

OpenStudy (amistre64):

www.wolframalpha.com

OpenStudy (amistre64):

its a math calculator that does all sorts of neat things

OpenStudy (anonymous):

oh I tried that and it was giving me the def. of find

OpenStudy (anonymous):

Lol i use my Chemistry one my boss gave me :P

OpenStudy (anonymous):

what is it ?

OpenStudy (amistre64):

then we can take that derivative part and plug in /5

OpenStudy (amistre64):

.5

OpenStudy (anonymous):

wait I look at the first result is that it ? (84 x^16+2)\/x^3

OpenStudy (amistre64):

yes, simplified of course

OpenStudy (anonymous):

ahh you made it look so easy

OpenStudy (amistre64):

:)